Suppose that replacement times for refrigerators are normally distributed with a mean of 9.7 years and a standard deviation of 1.1 years.
a) Find the replacement time that separates the bottom 18% from the top 82%.
Answer #1

Given that,

mean = = 9.7

standard deviation = = 1.1

Using standard normal table,

P(Z < z) = 18%

= P(Z < z) = 0.18  

= P(Z <-0.92 ) = 0.18

z = -0.92 Using standard normal z table,

Using z-score formula  

x= z * +

x= -0.92*1.1+9.7

x= 8.688